Footsteps Co. has a bond outstanding with a coupon rate of 5.4 percent and annual payments. The bond currently sells for $1,007.49, matures in 18 years, and has a par value of $1,000. What is the YTM of the bond?
Answer #1

use financial calculator as below

FV=1000

N=18

PMT=1000*5.4%=54

PV=-1007.49

Click CPT

Click I/Y=5.33% IS YTM

the above is answer..